# Cryoscope: A Cryogenic Infrared Survey Telescope

### v1

@@ 3.2 Critical objectives
Pre-shipment testing of the entire system is underway with the expected deployment date for Dome C in December 2025. The critical mission objectives for the Pathfinder are the following:
@@
    Validate image quality and focus control of the fully assembled cryogenic, double meniscus optical design. The next steps are to show that the optical performance revealed in the lab can be replicated on-sky when cold. On-sky image quality will be verified at Caltech ahead of deployment to Dome C for the 2025 season. Final validation will require that the optical performance not be degraded by atmospheric pressure on the window or cryogenic cooling of the primary mirror and field flatteners. Cold tests will demonstrate focus control and stability, accomplished by adjusting the operating temperature within the cryostat. Temperature variations of 1 C correspond to 10 $\,\mu$m of defocus. Shims are used to set the nominal focus position at a specific operating temperature. The adjustment range via thermal control will be validated, assuring that once shimmed correctly, no further adjustments will be required.
@@
    Suppress the telescope’s thermal self-emission below Dome C sky background levels. The novel double meniscus optical design enables one corrector lens to serve as a vacuum window so the telescope’s interior can be held at cryogenic temperatures. Suppression of instrumental thermal emission ensures that observations are sky-background-limited without limiting the field-of-view. Lab tests will demonstrate this concept of thermal background suppression by first capping the entrance aperture with a reflective mirror and measuring the background signal as a function of mirror temperature. Extrapolating the background signal at the mirror’s coldest temperature to the background expected in the absence of the mirror’s thermal emission, we will quantify the signal that arises exclusively from thermal self-emission. Following deployment at Dome C, the background from the telescope will be measured again using the reflective cap, but this time at much lower ambient operating temperature. After the emission by the telescope has been shown to be subdominant, Pathfinder will measure the K-dark sky brightness throughout the observing season.
@@ 3.3 Implementation at Concordia
Figure 12: Left: ASTEP telescope and mount in operation (Photo: V. Deloupy, 14 June 2024). Middle: CAD design of Cryoscope Pathfinder on the twin ASTEP direct drive mount. Right: Installation by the IPEV and PNRA teams at the Concordia Station of the pillar and dome to be used by Cryoscope Pathfinder, at latitude 75.100S, longitude 123.319E, and an elevation of 3233 m. (Photo: A. Agabi, 21 Nov. 2024).
@@
In November and December 2024, the IPEV and PNRA teams at Concordia installed a pillar, dome, and conduits designed to house power and data cables, as well as the necessary infrastructure to operate Cryoscope Pathfinder in December 2025.
@@ 4 Science cases
Next, we describe eight science cases that drive the Cryoscope design. While this list is not exhaustive, it illustrates the breadth in astronomy enabled by a wide-field infrared surveyor.
@@ 4.1 Exoplanets
Exoplanetology is a rapidly evolving field, with unexpected discoveries occurring routinely. This science case benefits from a robust fleet of operational and planned spacecraft. For the detection of new exoplanets, NASA’s TESS (Transiting Exoplanet Survey Satellite; Ricker et al., 2014) continues to deliver thousands of candidates, driving extensive follow-up work involving additional photometry, radial velocity measurements, and direct imaging campaigns to confirm and characterize these discoveries. In parallel, the focus of the field has increasingly shifted towards atmospheric characterization. Pioneered by the Hubble Space Telescope, this effort has entered a transformative phase with the advent of the James Webb Space Telescope (JWST), launched in 2021, which provides unprecedented high signal-to-noise spectra for robust atmospheric studies (e.g., Ahrer et al., 2023).
@@
Looking for temperate planets: Discovering long-period planets, particularly those in temperate zones conducive to habitability, is a critical goal for exoplanet science. These planets, often located far from their stars, provide essential clues about the diversity of planetary systems and their potential for hosting life. However, their detection remains challenging due to their faint signals and the extended time required to observe their orbits. Ground-based observatories, such as the ASTEP (Antarctic Search for Transiting Planets) program (Guillot et al., 2015; Crouzet et al., 2020; Schmider et al., 2022; Dransfield et al., 2022), have demonstrated the potential of Antarctica’s unique conditions – long, uninterrupted nights, and excellent visibility of the continuous viewing zones of key space missions like TESS, JWST, and soon Ariel.
@@
Figure 13: TESS planetary candidates observed by ASTEP from 2022 to 2024, plotted by K-band Vega magnitude versus orbital period. ASTEP’s focus on long-period planets has led to the discovery of several transiting planets with orbital periods exceeding 100 days. All these planetary transits would be observable with high signal-to-noise ratio using Cryoscope, and over 20% of them (highlighted) would be detectable with Cryoscope Pathfinder.
@@
Modest telescopes have already shown their potential. SPECULOOS excels in studying ultra-cool dwarfs due to its combination of aperture size and red sensitivity (Delrez et al., 2022), and UltraCAM is optimal for white dwarfs thanks to its fast readout times (Dhillon et al., 2007). By focusing on the K-dark band, Cryoscope aims to make groundbreaking observations, with the Pathfinder playing a critical role in ensuring we meet these ambitious goals.
@@ 4.2 Neutron Star Mergers
On August 17, 2017, the groundbreaking discovery of both gravitational waves and electromagnetic radiation from a binary neutron star (NS) merger marked the dawn of a new era in multi-messenger astrophysics (Abbott et al., 2017a). GW170817 lit up the entire electromagnetic spectrum, spanning $\gamma$-rays to radio, and yielded a scientific bonanza in fields as wide-ranging as strong field gravity, nucleosynthesis, extreme states of nuclear matter, the astrophysics of relativistic explosions, and even cosmology (Abbott et al., 2017b). For the first time, we saw direct evidence of r-process nucleosynthesis, the process by which half of the elements in the periodic table heavier than iron are synthesized (Lattimer & Schramm, 1976). Heavy line blanketing from the large density of bound-bound transitions renders the opacity of r-process rich matter much higher than the conventional iron peak elements, shifting the emergent spectrum of the “kilonova” out of the optical bands and into the infrared (Metzger et al., 2010; Barnes & Kasen, 2013; Kasen et al., 2013). Therefore, the infrared data are key to understanding the nucleosynthesis. Analysis of the infrared photometric evolution and the vivid broad features in the infrared spectroscopic sequence showed that at least 0.05 solar masses of heavy elements were synthesized by GW170817 (Coulter et al., 2017; Drout et al., 2017; Evans et al., 2017; Kasliwal et al., 2017; Smartt et al., 2017; Soares-Santos et al., 2017; Cowperthwaite et al., 2017; Arcavi, 2018). Combining the observed ejecta mass with rate estimates indeed gives numbers in the ballpark of the observed r-process abundances in the solar neighborhood. However, it is much debated whether the distribution of elemental abundances resembles the solar distribution. Evidence for the synthesis of the heaviest elements comes from late-time mid-infrared studies (Kasliwal et al., 2019).
@@
To answer these questions by localizing additional gravitational wave events, many wide-field optical cameras are being developed at all scales, including the Rubin Observatory. However, possibly none of the NS-BH mergers and only a subset of NS-NS mergers may radiate in the optical (Kyutoku et al., 2013; Kasen et al., 2017; Bulla, 2023). Specifically, it is predicted that the optical emission is limited to polar viewing angles in NS-NS mergers where the velocities are relatively high or the opacities are relatively low due to neutrino irradiation (Tanaka et al., 2014). If the merger remnant collapses promptly to form a BH, the optical emission is suppressed (Foucart et al., 2018). Despite extensive searches (e.g., with GROWTH (Kasliwal et al., 2020), GRANDMA (Antier et al., 2020), GOTO (Gompertz et al., 2020), J-GEM collaboration (Sasada et al., 2021)), no kilonova was identified during the LIGO-Virgo third observing run (April 1, 2019 to March 27, 2020) and first part of the fourth observing run (Ahumada et al., 2024, Pillas et al., in prep). This further constrains the optical luminosity function and we can now infer that the majority of kilonovae are not as intrinsically luminous as GW170817 (Kasliwal et al., 2020).
@@
Figure 14: Multi-band photometry of GW170817/AT2017gfo from Kasliwal et al. (2017). GW170817/AT2017gfo faded rapidly in the blue bands but was long-lived in the infrared. No K-dark data was obtained. Extrapolating from Ks using Cryoscope’s K-dark filter and the best-fit kilonova model from Kasen et al. (2017), the kilonova was brighter than -15 mag for over 5 days.
Figure 15: Model grids for kilonovae ejecta mass, velocity and lanthanide fraction (Kasen et al., 2017). Cryoscope will be sensitive to detecting kilonovae spanning a wide range of parameter space — all blue dots indicate the kilonova is brighter than -15 mag for more than two days (size of the symbol proportional to duration). NS-BH mergers are expected to have a larger lanthanide fraction than NS-NS mergers like GW170817.
Figure 16: Model for kilonova evolution as a function of time. Overplotted is the optical efficiencies in Rubin’s filters vs. Cryoscope’s K-dark.
Examining GW170817, we find that the light curve evolution was fastest in the blue filters and slowest in the red filters (Figure 14). Specifically, in K-dark, the kilonova stays brighter than -15 mag for over 5 days. Furthermore, if we examine kilonova model grids which depend on ejecta mass, ejecta velocity and lanthanide fraction, we find that a depth of -15 mag would be sensitive to most of the model phase space (Figure 15). Enhanced optical efficiency in K-dark at Dome C guarantees that Cryoscope is well-suited to detect and monitor the luminosity evolution of the most lanthanide-rich merger scenarios (Figure 16).
Figure 17: Simulation of neutron star mergers for the fifth gravitational wave observing run planned at the end of this decade. Color-coding indicates the exposure time needed for Cryoscope to map the 90% GW localization area to find a kilonova brighter than -15 mag. Light red points indicate events in the Northern hemisphere that are not accessible from the Antarctic location. Note that Cryoscope can detect all NS mergers within 300 Mpc irrespective of the localization and even detect well-localized NS mergers out to 1 Gpc.
@@ 4.3 Galactic Stellar Science
Cryoscope and its fifth-scale Pathfinder will enable many facets of Galactic stellar science. Here, we present a few exemplars.
Stellar mergers, particularly close binary systems undergoing mass transfer and subsequent merger, have a profound impact on stellar mass function (Wang et al., 2022). These mergers give rise to short period binary systems (Tauris & van den Heuvel, 2006) and may be observed as luminous outbursts of the primary star followed by the slow ejection of the cool outer envelope and the formation of dust (Ivanova et al., 2013). Understanding them is crucial for studying stellar populations, but their identification is challenging due to their predominantly infrared emission as shown in Figure 18 (MacLeod et al., 2022). Cryoscope Pathfinder aims to overcome this challenge by efficiently detecting stellar mergers through a combination of short-lived optical outbursts and long-lived infrared emission. It provides valuable insights into the demographics of these mergers, including the elusive lowest mass mergers that are difficult to identify through optical surveys.
Figure 18: Evolutionary tracks (MacLeod et al., 2022) for merging binaries with different primary masses color-coded to denote the expected extinction (see color-bar on the right). Upper limits are shown for the merger evolutionary stages accessible to the Rubin Observatory $r$-band (light blue) and for Cryoscope Pathfinder (red) for a Galactic merger at 10 kpc. The Pathfinder sensitivity rivals that of Rubin due to the K- dark bandpass being much more immune to extinction than the visible bandpass. The expected rate of discovery is about one per year in the Milky Way and up to a few per year in the nearby galaxies. Even a single such discovery by Cryoscope Pathfinder would likely lead to a high-impact scientific result as it probes a hitherto unexplored region of parameter space in stellar merger physics. If the Pathfinder does not see such extremely extincted mergers, it would severely constrain theoretical models.
@@
Ultracompact binaries (UCBs) are important to study accretion physics in a unique regime. UCBs with white dwarf components emit both electromagnetic and gravitational radiation, enabling a better characterization of the source properties. Understanding UCB formation is crucial for studying Type Ia supernova progenitors and binary evolution. Optical surveys have been instrumental, but detecting UCBs with cold companions (as a subclass called AM CVns $\sim 10$), is challenging. Cryoscope Pathfinder’s K-band search can overcome the shortcomings in the optical, discovering tens of non-outbursting AM CVns within a year. This will enhance our understanding of nearby gravitational wave emitters to prepare for the LISA space mission (Breivik et al., 2018; van Roestel et al., 2021).
@@
Young stellar objects: Star formation occurs through the collapse of molecular cloud cores (see Armitage, 2019, for a review). This process involves mass transfer from the cloud to the central protostar, marked by notable accretion outbursts. However, a comprehensive understanding of the amplitudes and timescales of these outbursts is still lacking. The challenge lies in the fact that most of the process is invisible in optical bands, where current time domain surveys excel in detecting outbursts. In contrast, infrared bands offer an opportunity to characterize these outbursts throughout all phases of the accretion process (Armitage, 2019; Fischer et al., 2023). The Cryoscope Pathfinder’s time-domain survey of the Galactic plane will systematically uncover hundreds of young star outbursts in the visible Galactic disk, while complementary follow-up will complete the observations.
@@
Novae: Classical novae constitute some of the oldest known explosive transients, caused by thermonuclear ignition on the surface of an accreting white dwarf (Starrfield et al., 2016). In a small fraction $\sim 10\%$ of cases, classical novae form dust shortly (weeks to months) after outburst, as characterized by dramatic dips in their optical light curves (Williams et al., 2013). Recent work suggests an intricate link between dust formation and their identification as luminous $\gamma$-ray transients (Chomiuk et al., 2021). A theoretical prediction that dust formation should be ubiquitous in novae, but observable as optical dips only along lines of sight near the binary orbital plane (Derdzinski et al., 2017) results in a testable prediction. As some of the brightest transients in the infrared sky, Galactic novae thus provide the ideal opportunity to test these predictions in K-band light curves. Cryoscope Pathfinder will be sensitive to all Galactic novae in the southern sky, where dust formation can be easily detected, and will reveal a large number of systems that currently elude detection in optical wavelengths (De et al., 2021). When combined with multi-wavelength data from $\gamma$-ray and radio observatories, these observations can reveal the physics of shocks in these explosions and shed light on the nature of the most luminous shock-powered explosions in the universe such as superluminous supernovae.
@@
RR Lyrae: RR Lyrae (RRL) variables are believed to be among the oldest stars in the Galaxy, so are important probes for Galactic archaeology. They are easy to recognize by their characteristic light curves. While most RRL stars near the Sun are old ($>10$ Gyr) and metal-poor ([Fe/H] $<-1$) and are part of the Galactic stellar halo, there is a sub-population of metal-rich RRL stars in the solar neighborhood with abundances [Fe/H] $>-1$ that are clearly part of the Galactic disk. They rotate rapidly around the Galaxy and their velocity dispersion is low. Recent studies of a sample of a few hundred RRL stars with accurate abundances and kinematics show that these apparently old, metal-rich RRL stars have kinematics similar to relatively young stars of the Galactic thin disk. The origin and evolution of these metal-rich disk RRL stars is not understood. RRL stars of similar [Fe/H] are found in the inner Galaxy, and it is possible that their counterparts in the solar neighborhood have migrated from the inner Galaxy via radial migration. The migration process can generate a kinematically cold population in the outer disk (e.g. Vera-Ciro et al., 2014). The radial migration process itself is not well understood. The goal of this science is to expand our understanding of (i) the metal-rich disk RRL stars near the Sun, and (ii) how radial migration from the inner Galaxy contributes to the stellar populations of the outer disk.
A kinematically cold population of RRL stars in the outer Galactic disk is likely to have a strong concentration to the Galactic plane. So far, there is little data available on RRL star distributions in the low Galactic latitude disk where the interstellar extinction is high. The Cryoscope survey telescope, observing in the K-dark band, would be able to search for these expected low-latitude RRL stars. Such a search can be conducted by observing two Pathfinder fields centered at galactic latitude $b=0$ and galactic longitude $l$ in the range 280 to 320. Each field would be observed about 30 times over about a year with irregular cadence. The exposure times for each observations would be long enough to reach a KAB mag of 15, with errors of about 2 percent. The periods of metal-rich RRL stars are mostly in the range 0.35 to 0.5 days.

### v2

@@ 3.2 Critical objectives
Pre-shipment testing of the entire system is underway with the expected deployment date for Dome C in December 2026. The critical mission objectives for the Pathfinder are the following:
@@
    Validate image quality and focus control of the fully assembled cryogenic, double meniscus optical design. The next steps are to show that the optical performance revealed in the lab can be replicated on-sky when cold. On-sky image quality will be verified at Caltech ahead of deployment to Dome C for the 2026 season. Final validation will require that the optical performance not be degraded by atmospheric pressure on the window or cryogenic cooling of the primary mirror and field flatteners. Cold tests will demonstrate focus control and stability, accomplished by adjusting the operating temperature within the cryostat. Temperature variations of 1 C correspond to 10 $\,\mu$m of defocus. Shims are used to set the nominal focus position at a specific operating temperature. The adjustment range via thermal control will be validated, assuring that once shimmed correctly, no further adjustments will be required.
@@
    Suppress the telescope’s thermal self-emission below Dome C sky background levels. The novel double meniscus optical design enables one corrector lens to serve as a vacuum window so the telescope’s interior can be held at cryogenic temperatures. Suppression of instrumental thermal emission ensures that observations are sky-background-limited without limiting the field-of-view. Lab tests will demonstrate this concept of thermal background suppression by first capping the entrance aperture with a reflective mirror and measuring the background signal as a function of mirror temperature. Extrapolating the background signal at the mirror’s coldest temperature to the background expected in the absence of the mirror’s thermal emission, we will quantify the signal that arises exclusively from thermal self-emission. Following deployment at Dome C, the background from the telescope will be measured again using the reflective cap, but this time at much lower ambient operating temperature. After the emission by the telescope has been shown to be subdominant, Pathfinder will measure the $K_{dark}$ sky brightness throughout the observing season.
@@ 3.3 Implementation at Concordia
Figure 11: Left: ASTEP telescope and mount in operation (Photo: V. Deloupy, 14 June 2024). Middle: CAD design of Cryoscope Pathfinder on the twin ASTEP direct drive mount. The Pathfinder will be installed on a mount identical to the one used by ASTEP. Right: Installation by the IPEV and PNRA teams at the Concordia Station of the pillar and dome to be used by Cryoscope Pathfinder, at latitude 75.100S, longitude 123.319E, and an elevation of 3233 m. (Photo: A. Agabi, 21 Nov. 2024).
@@
In November and December 2024, the IPEV and PNRA teams at Concordia installed a pillar, dome, and conduits designed to house power and data cables, as well as the necessary infrastructure to operate Cryoscope Pathfinder in December 2026.
Concordia Station is connected to the Internet via a satellite link. Although the available bandwidth is limited, it provides a data transmission capacity of 20 GB/day for Cryoscope Pathfinder. Until bandwidth capacity is increased, data processing will be performed using pipelines deployed on an on-site server, with results subsequently transmitted. The station’s Internet connection enables remote control of the telescope’s operation and data acquisition.
@@ 4 Science cases
Next, we describe nine science cases that drive the Cryoscope design. While this list is not exhaustive, it illustrates the breadth in astronomy enabled by a wide-field infrared surveyor.
@@ 4.1 Exoplanets
Exoplanetology is a rapidly evolving field, with unexpected discoveries occurring routinely. This science case benefits from a robust fleet of operational and planned spacecraft. For the detection of new exoplanets, NASA’s TESS (Transiting Exoplanet Survey Satellite; Ricker et al., 2014) has produced over one thousand candidates in the past two years, outpacing, and also driving an extensive follow-up work involving additional photometry, radial velocity measurements, and direct imaging campaigns to confirm and characterize these discoveries. In parallel, the focus of the field has increasingly shifted towards atmospheric characterization. Pioneered by the Hubble Space Telescope, this effort has entered a transformative phase with the advent of the James Webb Space Telescope ( JWST), launched in 2021, which provides unprecedented high signal-to-noise spectra for robust atmospheric studies (e.g., Ahrer et al., 2023).
@@
Looking for temperate planets: Discovering long-period planets, particularly those in temperate zones conducive to habitability, is a critical goal for exoplanet science. These planets, often located far from their stars, provide essential clues about the diversity of planetary systems and their potential for hosting life. However, their detection remains challenging due to their faint signals and the extended time required to constrain their orbits. Ground-based observatories, such as the ASTEP (Antarctic Search for Transiting Planets) program (Guillot et al., 2015; Crouzet et al., 2020; Schmider et al., 2022; Dransfield et al., 2022), have demonstrated the potential of Antarctica’s unique conditions – long, uninterrupted nights, and excellent visibility of the continuous viewing zones of key space missions like TESS, JWST, and soon Ariel.
@@
Figure 12: TESS planetary candidates observed by ASTEP from 2022 to 2024, plotted by $K$-band Vega magnitude versus orbital period. ASTEP’s focus on long-period planets has led to the discovery of several transiting planets with orbital periods exceeding 100 days. All these planetary transits would be observable with high signal-to-noise ratio using Cryoscope, and over 20% of them (highlighted) would be detectable with Cryoscope Pathfinder. The color bar illustrates the transit depth produced by an exoplanet, in ppt (parts per thousand in relative photometry).
@@
Modest telescopes have already shown their potential. SPECULOOS excels in studying ultra-cool dwarfs due to its combination of aperture size and red sensitivity (Delrez et al., 2022), and UltraCAM is optimal for white dwarfs thanks to its fast readout times (Dhillon et al., 2007). By focusing on the $K_{dark}$ band, Cryoscope aims to make groundbreaking observations, with the Pathfinder playing a critical role in ensuring we meet these ambitious goals.
@@ 4.2 Neutron Star Mergers
On August 17, 2017, the groundbreaking discovery of both gravitational waves and electromagnetic radiation from a binary neutron star (NS) merger marked the dawn of a new era in multi-messenger astrophysics (Abbott et al., 2017a). GW170817 lit up the entire electromagnetic spectrum, spanning $\gamma$-rays to radio, and yielded a scientific bonanza in fields as wide-ranging as strong field gravity, nucleosynthesis, extreme states of nuclear matter, the astrophysics of relativistic explosions, and even cosmology (Abbott et al., 2017b). For the first time, we saw direct evidence of r-process nucleosynthesis, the process by which half of the elements in the periodic table heavier than iron are synthesized (Lattimer & Schramm, 1976). Heavy line blanketing from the large density of bound-bound transitions renders the opacity of r-process rich matter much higher than the conventional iron peak elements, shifting the emergent spectrum of the “kilonova” out of the optical bands and into the infrared (Metzger et al., 2010; Barnes & Kasen, 2013; Kasen et al., 2013).
The infrared data are therefore key to understanding the nucleosynthesis of heavy elements by the r-process. Analysis of the infrared photometric evolution and the vivid broad features in the infrared spectroscopic sequence showed that at least 0.05 solar masses of heavy elements were synthesized by GW170817 (Coulter et al., 2017; Drout et al., 2017; Evans et al., 2017; Kasliwal et al., 2017; Smartt et al., 2017; Soares-Santos et al., 2017; Cowperthwaite et al., 2017; Arcavi, 2018). Combining the observed ejecta mass with rate estimates indeed gives numbers in the ballpark of the observed r-process abundances in the solar neighborhood. However, it is much debated whether the distribution of elemental abundances resembles the solar distribution. Evidence for the synthesis of the heaviest elements comes from late-time mid-infrared studies (Kasliwal et al., 2019).
@@
To answer these questions by localizing additional gravitational wave events, many wide-field optical cameras are being developed at all scales, including the Rubin Observatory. However, possibly none of the NS-BH mergers and only a subset of NS-NS mergers may radiate in the optical (Kyutoku et al., 2013; Kasen et al., 2017; Bulla, 2023). Specifically, it is predicted that the optical emission is limited to polar viewing angles in NS-NS mergers where the velocities are relatively high or the opacities are relatively low due to neutrino irradiation (Tanaka et al., 2014). If the merger remnant collapses promptly to form a BH, the optical emission is suppressed (Foucart et al., 2018). Despite extensive searches (e.g., with GROWTH (Kasliwal et al., 2020), GRANDMA (Antier et al., 2020), GOTO (Gompertz et al., 2020), J-GEM collaboration (Sasada et al., 2021)), no kilonova was identified during the LIGO-Virgo third observing run (April 1, 2019 to March 27, 2020) and first part of the fourth observing run (Ahumada et al., 2024, Pillas et al., in prep). The small number of BNS mergers in the third and fourth observing runs have lowered their volumetric rate estimate. Furthermore, no additional kilonova detection since GW170817 also constrains the optical luminosity function and we can now infer that the majority of kilonovae are not as intrinsically luminous in the optical as GW170817 (Kasliwal et al., 2020; Ahumada et al., 2024).
@@
Figure 13: Multi-band photometry of GW170817/AT2017gfo from Kasliwal et al. (2017). GW170817/AT2017gfo faded rapidly in the blue bands but was long-lived in the infrared. No $K_{dark}$ data was obtained. Extrapolating from $K_{s}$ using Cryoscope’s $K_{dark}$ filter and the best-fit kilonova model from Kasen et al. (2017), the kilonova was brighter than $-15$  mag for over 5  days.
Figure 14: Model grids for kilonovae ejecta mass, velocity and lanthanide fraction (Kasen et al., 2017). Cryoscope will be sensitive to detecting kilonovae spanning a wide range of parameter space. All blue dots indicate the kilonova is brighter than $-15$  mag for more than two days (size of the symbol proportional to duration). NS-BH mergers are expected to have a larger lanthanide fraction than NS-NS mergers like GW170817.
Figure 15: Three model kilonovae spectra shown at three snapshots in time after the explosion. The expected total optical throughput (right hand vertical axis) in the six Rubin filters is also shown, for comparison to that of Cryoscope’s $K_{dark}$.
Examining GW170817, we find that the light curve evolution was fastest in the blue filters and slowest in the red filters (Figure 13). Specifically, in $K_{dark}$, the kilonova stays brighter than $-15$  mag for over 5  days. Furthermore, if we examine kilonova model grids which depend on ejecta mass, ejecta velocity and lanthanide fraction, we find that a depth of $-15$  mag would be sensitive to most of the model phase space (Figure 14). Enhanced optical efficiency in $K_{dark}$ at Dome C guarantees that Cryoscope is well-suited to detect and monitor the luminosity evolution of the most lanthanide-rich merger scenarios (Figure 15).
Figure 16: Simulation of neutron star mergers for the fifth gravitational wave observing run planned at the end of this decade. Color-coding indicates the exposure time needed for Cryoscope to map the 90% GW localization area to find a kilonova brighter than $-15$  mag. Gray points indicate events that are undetectable, requiring more than 105 seconds to tile the region to required depths. Light red points indicate events in the Northern hemisphere that are not accessible from the Antarctic location. Note that Cryoscope can detect all NS mergers within 300 Mpc irrespective of the localization and even detect well-localized NS mergers ($<300$  deg2 credible regions) out to 1 Gpc.
@@ 4.3 Galactic Stellar Science
Cryoscope and its fifth-scale Pathfinder will enable many facets of Galactic stellar science. Here, we present a few examples.
Stellar mergers, particularly close binary systems undergoing mass transfer and subsequent merger, have a profound impact on stellar mass function (Wang et al., 2022). These mergers give rise to short period binary systems (Tauris & van den Heuvel, 2006) and may be observed as luminous outbursts of the primary star followed by the slow ejection of the cool outer envelope and the formation of dust (Ivanova et al., 2013). Understanding them is crucial for studying stellar populations, but their identification is challenging due to their predominantly infrared emission as shown in Figure 17 (MacLeod et al., 2022). Cryoscope Pathfinder aims to overcome this challenge by efficiently detecting stellar mergers through a combination of short-lived optical outbursts and long-lived infrared emission. It provides valuable insights into the demographics of these mergers, including the elusive lowest-mass mergers, which are difficult to identify through optical surveys.
Figure 17: Evolutionary tracks (MacLeod et al., 2022) for merging binaries with different primary masses color-coded to denote the expected extinction (see color-bar on the right). Upper limits are shown for the merger evolutionary stages accessible to the Rubin Observatory $r$-band (light blue) and for Cryoscope Pathfinder (red) for a Galactic merger at 10 kpc. The Pathfinder sensitivity rivals that of Rubin due to the $K_{dark}$ bandpass being much more immune to extinction than the visible bandpass. The expected rate of discovery is about one per year in the Milky Way and up to a few per year in the nearby galaxies. Even a single such discovery by Cryoscope Pathfinder would likely lead to a high-impact scientific result as it probes a hitherto unexplored region of parameter space in stellar merger physics. If the Pathfinder does not see such extremely extincted mergers, it would severely constrain theoretical models.
@@
Ultracompact binaries (UCBs) are important for characterizing accretion physics in a unique regime. UCBs with white dwarf components emit both electromagnetic and gravitational radiation, enabling a better characterization of the source properties. Understanding UCB formation is crucial for studying Type Ia supernova progenitors and binary evolution. Detecting UCBs with cold companions (as a subclass called AM CVns $\sim 10$), is challenging. Cryoscope Pathfinder’s $K$-band search can overcome the shortcomings of visible-wavelength surveys and reveal tens of non-outbursting AM CVns within a year. This will enhance our understanding of nearby gravitational wave emitters to prepare for the LISA space mission (Breivik et al., 2018; van Roestel et al., 2021).
@@
Young stellar objects: Star formation occurs through the collapse of molecular cloud cores (see Armitage, 2019, for a review). This process involves mass transfer from the cloud to the central protostar, marked by notable accretion outbursts. However, a comprehensive understanding of the amplitudes and timescales of these outbursts is still lacking. The challenge lies in the fact that most of the process is invisible in optical bands, where current time domain surveys excel in detecting outbursts. In contrast, infrared bands offer an opportunity to characterize these outbursts throughout all phases of the accretion process (Armitage, 2019; Fischer et al., 2023). The Cryoscope Pathfinder’s time-domain survey of the Galactic plane will systematically uncover hundreds of young star outbursts in the visible Galactic disk.
@@
Novae: Classical novae constitute some of the oldest known explosive transients, caused by thermonuclear ignition on the surface of an accreting white dwarf (Starrfield et al., 2016). In a small fraction ($\sim 10\%$) of cases, classical novae form dust shortly (weeks to months) after outburst, as characterized by dramatic dips in their optical light curves (Williams et al., 2013). Recent work suggests an intricate link between dust formation and their identification as luminous $\gamma$-ray transients (Chomiuk et al., 2021). A theoretical prediction that dust formation should be ubiquitous in novae, but observable as optical dips only along lines of sight near the binary orbital plane (Derdzinski et al., 2017) results in a testable prediction. As some of the brightest transients in the infrared sky, Galactic novae thus provide the ideal opportunity to test these predictions in $K$-band light curves. Cryoscope Pathfinder will be sensitive to all Galactic novae in the southern sky, where dust formation can be easily detected, and will reveal a large number of systems that currently elude detection in optical wavelengths (De et al., 2021). When combined with multi-wavelength data from $\gamma$-ray and radio observatories, these observations can reveal the physics of shocks in these explosions and shed light on the nature of the most luminous shock-powered explosions in the universe such as superluminous supernovae.
@@
RR Lyrae: RR Lyrae (RRL) variables are believed to be among the oldest stars in the Galaxy, so are important probes for Galactic archaeology. They are easy to recognize by their characteristic light curves. While most RRL stars near the Sun are old ($>10$ Gyr) and metal-poor ([Fe/H] $<-1$) and are part of the Galactic stellar halo, there is a sub-population of metal-rich RRL stars in the solar neighborhood with abundances [Fe/H] $>-1$ that are clearly part of the Galactic disk. They rotate rapidly around the Galaxy and their velocity dispersion is low. Recent studies of a sample of a few hundred RRL stars with accurate abundances and kinematics show that these apparently old, metal-rich RRL stars have kinematics similar to relatively young stars of the Galactic thin disk. The origin and evolution of these metal-rich disk RRL stars is not understood. RRL stars of similar [Fe/H] are found in the inner Galaxy, and it is possible that their counterparts in the solar neighborhood have migrated from the inner Galaxy via radial migration. The migration process can generate a kinematically cold population in the outer disk (e.g. Vera-Ciro et al., 2014). The radial migration process itself is not well understood. With Cryoscope Pathfinder observations, we will expand our understanding of (i) the metal-rich disk RRL stars near the Sun, and (ii) how radial migration from the inner Galaxy contributes to the stellar populations of the outer disk.
A kinematically cold population of RRL stars in the outer Galactic disk is likely to have a strong concentration to the Galactic plane. So far, there is little data available on RRL star distributions in the low Galactic latitude disk where the interstellar extinction is high. The Cryoscope survey telescope, observing in the $K_{dark}$ band, would be able to search for these expected low-latitude RRL stars. Such a search can be conducted by observing two Pathfinder fields centered at galactic latitude $b=0$ and galactic longitude $l$ in the range 280 to 320. Each field would be observed about 30 times over about a year with irregular cadence. The exposure times for each observation would be long enough to reach a $K_{\rm AB}$  mag of 15, with errors of about 2 percent. The periods of metal-rich RRL stars are mostly in the range 0.35 to 0.5 days.